Grade 11Magical Mathematics[Interesting Approach]

if a,b,c are in A.P ,Then prove that b+c-a,c+a-b,a+b-c are also in A.P

talha khan
14 Years agoGrade 11

1 Answer

Vaishal Shah
ApprovedApproved Tutor Answer14 Years ago

2b=a+c

Now,

2(c+a-b)=b+c-a+a+b-c

2c+2a-2b=2b

2c+2a=4b

c+a=2b